HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

SELF-BROADENING COEFFICIENTS FOR FUNDAMENTAL AND HOT BANDS OF ACETYLENE IN THE 1300 CM^{−1} REGION

\underline{N. DRICOT} , J. CLÉMENT, O. BROWET, B. VISPOEL, M. LEPÈRE , Research unit Lasers and Spectroscopies (LLS), Institute of Life, Earth and

Environment (ILEE), University of Namur, 61, Rue de Bruxelles, Namur, Belgium

Acetylene is one of the most common hydrocarbons present in Eath’s atmosphere^{1} which origiantes mainly from human activities^{2}, but also from natural phenomena^{3}. Because of its presence in combustion processes, it can be used as marker for antropogenic pollution monitoring. C_{2}H_{2} was also detected in planetary atmospheres inside and outside our solar system, as well as in interstellar medium (see review in Ref^{4}). The computation of transfer radiative models and opacities requires accurate spectroscopic data. In particular, the line shape parameters are the least well known.

In this study, the self-broadening coefficients were determined for acetylene lines in the 1300 cm^{−1} spectral region. The measurements were performed at room temperature using a high-resolution mid-infrared dual comb spectrometer based on quantum cascade lasers, that is well suited for line shape parameter measurements^{5}. Thanks to a specifically designed absorption cell, transitions in the ν_{4} + ν_{5} fundamental and two hot bands were studied. The obtained results exhibit the expected smooth variation with respect to the rotational quantum number J. To the best of our knowledge, it is the first time that the self-broadening coefficients were directly determined for the two studied hot bands. The results for the fundamental band are in good agreement with literature.

Acknowledgements: The authors thank the IRsweep company for the support during the measurements performed with the dual comb spectrometer.

1doi:10.1007/s00340-010-4027-5F.M. Schmidt et al., ”Direct detection of acetylene in air by continuous wave cavity ring-down spectroscopy”, Applied Physics B Lasers and Optics, 101, 671–682 (2010)

2doi:10.1016/S1352-2310(96)00266-XA.M. Mohan Rao et al., ”Non-methane hydrocarbons in industrial locations of Bombay”, Atmospheric Environment, 31(7), 1077–1085 (1997)

3doi:10.1038/333051a0M. Kanakidou et al., ”Marine source of atmospheric acetylene”, Nature, 333, 51–52 (1988)

4doi:10.1093/mnras/staa229K.L. Chubb et al., ”ExoMol molecular line lists - XXXVII. Spectra of acetylene”, Monthly Notices of the Royal Astronomical Society, 493(2), 1531–1545 (2020)

5doi:10.1016/j.jqsrt.2022.108239M. Lepère et al., ”A mid-infrared dual-comb spectrometer in stepsweep mode for high-resolution molecular spectroscopy”, Journal of Quantitative Spectroscopy and Radiative Transfer, 287, 108239 (2022)

p-number: p231  Submitted on Wed, 21 Jun 2023 13:47:29 +0200